OpenStudy (anonymous):

In a search-and-rescue mission, a team maps out a circular search area from the last known location of a group of hikers. If the search region is 113.04 mi^2, how far from the last known location, to the nearest mile, is the team searching? Use pi=3.14

OpenStudy (anonymous):

area = 113.04 = pi*r^2 113.04 = 3.14 * r^2 r^2=113.04/3.14 r=sqrt(113.04/3.14) r=sqrt(36) r=6 So the search radius is 6 miles around the last known location of the hikers

OpenStudy (anonymous):

nice!! thnaks